Standardization of Sodium Thiosulfate

In one trial of the standardization of a Na2S2O3 solution, a 10.00 mL volume of 5.00×10-3 M KIO3 is pipetted into a 250 mL Erlenmeyer flask. The solution is titratedto the endpoint with 15.23 mL of the Na2S2O3 solution. What is the molar concentration of this sodium thiosulfate solution?
Answer #1

KIO3 + 6 Na2S2O3 + 6 H+ => KI + 3 Na2S4O6 + 6 Na+ + 3 H2O

Moles of KIO3 = volume x concentration of KIO3

= 10.00/1000 x 5.00 x 10-3 = 5.00 x 10-5 M

Moles of Na2S2O3 = 6 x moles of KIO3

= 6 x 5.00 x 10-5 = 3.00 x 10-4 mol

Concentration of Na2S2O3 = moles/volume in L of Na2S2O3

= 3.00 x 10-4/0.01523

= 0.0197 M = 1.97 x 10-2 M

answered by: Scout